# Co-rewarding: Stable Self-supervised RL for Eliciting Reasoning in Large Language Models

### v2

@@ 4.2.2 Further Analysis
Each part contributes to Co-rewarding. Table 5 summarizes the ablations. For Co-rewarding-I, replacing cross supervision with training only on original or rephrased data yields comparable results, but both underperform the full method. This indicates that the quality of original and rephrased data is similar, while the cross supervision across views provides the crucial gains. For Co-rewarding-II, removing the EMA update of the reference teacher model causes clear degradation, highlighting the necessity of teacher updates for improving pseudo-label quality.
@@
Case Study of the model reasoning with different learning methods. Figure 7 provides a concrete example to illustrate the qualitative difference between self-rewarding baselines and our Co-rewarding. Majority-Voting exhibits reward hacking by boxing an incorrect answer “0” to pursue consensus, even though the reasoning steps are correct. Entropy produces repetitive outputs as its decoding probability distribution collapses onto a narrow set of tokens during entropy minimization. In contrast, Co-rewarding generates coherent step-by-step reasoning and correctly boxes the final answer, showing its capacity to overcome reward hacking and elicit genuine reasoning.
@@ 5 Conclusion
In this work, we introduced Co-rewarding, a self-supervised RL framework that elicits the reasoning capability of LLMs through complementary supervision. Unlike prior self-rewarding methods that entangle rewards with single-view outputs and risk collapse, Co-rewarding establishes stability by decoupling the reward signal from the current online policy with the single-view output. Specifically, Co-rewarding-I leverages contrastive agreement across semantically analogous questions, while Co-rewarding-II employs a dynamically updated teacher to provide insulated pseudo-labels. Together, these designs construct cross-referable reward signals without explicit labels, aligning RL with invariances in reasoning rather than the mere correctness of isolated outputs. We hope this work will inspire further exploration into self-supervised RL for reasoning to advance the development.
@@ Appendix A Related Work
Large Language Model Reasoning. LLMs have shown impressive performance on vast tasks that require reasoning, including solving mathematical problems, writing code, and answering logical questions. One of the key techniques that has improved LLM reasoning is Chain-of-Thought (CoT) prompting (Wei et al., 2022). CoT encourages the model to generate intermediate reasoning steps before producing the final answer, which has been shown to enhance performance on tasks like arithmetic, commonsense reasoning, and symbolic reasoning. Subsequent work has extended CoT by integrating it with various strategies, including compositional generalization (Zhou et al., 2023a; Khot et al., 2023) and employing structural reasoning approaches (Yao et al., 2023a; Besta et al., 2024; Yang et al., 2024). In addition, CoT serves as a fundamental framework for techniques like fine-tuninig (Zelikman et al., 2022), argentic workflow (Yao et al., 2023b), and paving the way for inference-time scaling (Snell et al., 2024).
@@
RL for Large Language Models. Several RL algorithms have been developed primarily for alignment tasks. Specifically, DPO (Rafailov et al., 2023), CPO (Xu et al., 2024), and their variants (Li et al., 2024b; Guo et al., 2024; Munos et al., 2024; Hong et al., 2024; Xie et al., 2024) rely on pairs of outputs labeled by human preference. In contrast, KTO (Ethayarajh et al., 2024) and BCO (Jung et al., 2024) require only a single binary label (like or dislike) for each output. Besides, the PRM (Uesato et al., 2022; Lightman et al., 2024) and Step-KTO (Lin et al., 2025a) offer step-by-step guidance by incorporating feedback at each reasoning step rather than focusing solely on the final outputs. Recently, the follow-up work of GRPO improves the optimization objective, e.g., DAPO (Yu et al., 2025), Dr. GRPO (Liu et al., 2025a), REINFORCE++ (Hu, 2025), CPPO (Lin et al., 2025b), and GPG (Chu et al., 2025). Another line of research generalizes GRPO to broader applications such as multimodal reasoning (Zhou et al., 2025; Huang et al., 2025; Chu et al., 2025; Liu et al., 2025b; Zhang et al., 2025a) and logical reasoning (Xie et al., 2025).
@@
RL without External Reward. RL methods have shown promising scaling capabilities to enhance the reasoning abilities of LLMs (Guo et al., 2025), yet they are often limited by the availability of training data for reward signals (Gao et al., 2023; Liu et al., 2023). Notably, Wang et al. (Wang et al., 2025) demonstrate that RL can effectively bootstrap LLM reasoning with as little as a single training example, highlighting the potential to minimize or even eliminate reliance on external reward signals during training. Recent efforts leverage distinct strategies for reward assignment. For instance, SIRLC (Pang et al., 2024) and AZR (Zhao et al., 2025a) utilize an LLM-as-the-judge approach to assign rewards. In contrast, methods like SRT, TTRL, and their variants (Shafayat et al., 2025; Zuo et al., 2025; Fang et al., 2025; Zhang et al., 2025b) employ self-consistency (Wang et al., 2022) to generate pseudo-rewards, reducing dependence on external annotations. Meanwhile, INTUITOR, RLSC, and RENT (Zhao et al., 2025b; Li et al., 2025; Prabhudesai et al., 2025) harness the internal confidence scores of LLMs as intrinsic reward signals. Additionally, EMPO and its variants (Zhang et al., 2025c; Agarwal et al., 2025) promote reasoning by minimizing entropy during the reasoning process, further diversifying the approaches to incentivize robust LLM performance.
@@ Appendix B Pseudo Code of Co-rewarding
To intuitively present the pipeline of Co-rewarding, we summarize the pseudo codes of Co-rewarding-I and Co-rewarding-II in Algorithm 1 and Algorithm 2, respectively.
@@ 5. Output ONLY between the following markers, and strictly in this format (no extra explanation):
Settings Co-rewarding-I Co-rewarding-II
Batch Size 128 128
Max Prompt Length 512 512
Max Response Length 3072 3072
Train Steps 170-220 300-330
Learning Rate 3e-6 3e-6
# Policy Rollout $G$ 8 8
# Reference Rollout $\tilde{G} -$ 8
Clip Ratio 0.2 0.2
Warmup Style Cosine Cosine
Warmup Steps Ratio 0.1 0.1
KL Loss Coefficient 0.005 0.001
@@
Training Temperature 1.0 1.0
Evaluation Temperature 0.8 0.8
EMA $\alpha_{\textrm{start}}$ - 0.99
EMA $\alpha_{\textrm{end}}$ - 0.9999
@@
Open-RS (Dang & Ngo, 2025) 7,000 Training Set
@@ C.3 More Evaluation Details
Methods MATH500 GSM8K AMC LiveCode CRUX IFEval MMLU-Pro
@@
Before RL 53.6 19.48 10.69 9.95 18.50 29.83 32.50
- GT-Reward (Shao et al., 2024) 65.4 82.18 32.98 13.93 32.12 33.66 36.74
- Self-Certainty (Zhao et al., 2025b) 64.2 80.52 28.92 10.90 29.00 32.22 33.88
- Entropy (Prabhudesai et al., 2025) 63.2 80.44 29.67 9.05 29.00 32.94 35.35
- Majority-Voting (Shafayat et al., 2025) 64.6 82.41 33.13 14.03 36.38 35.19 35.50
- Co-rewarding-I (Ours) 65.4 84.53 30.57 16.40 36.88 33.86 36.38
- Co-rewarding-II (Ours) 65.2 81.72 32.38 22.25 40.25 32.74 30.79
@@
Before RL 69.4 24.71 15.81 3.79 26.38 38.19 44.76
- GT-Reward (Shao et al., 2024) 76.4 88.02 45.63 15.92 45.12 41.49 41.12
- Self-Certainty (Zhao et al., 2025b) 72.8 84.31 38.55 12.04 54.12 37.24 43.30
- Entropy (Prabhudesai et al., 2025) 72.2 81.43 39.61 16.49 51.88 40.33 42.79
- Majority-Voting (Shafayat et al., 2025) 74.4 84.53 40.96 15.45 51.00 38.60 43.35
- Co-rewarding-I (Ours) 74.6 89.61 41.27 15.73 55.58 42.86 40.51
- Co-rewarding-II (Ours) 73.6 89.31 42.77 8.25 47.50 41.82 37.45
@@
Before RL 57.0 19.56 8.43 4.45 7.50 33.65 33.00
- GT-Reward (Shao et al., 2024) 69.6 81.57 35.54 13.74 35.25 36.16 39.12
- Self-Certainty (Zhao et al., 2025b) 58.2 40.25 23.04 9.86 18.00 32.96 35.13
- Entropy (Prabhudesai et al., 2025) 63.6 71.79 31.63 13.74 31.37 35.37 36.67
- Majority-Voting (Shafayat et al., 2025) 65.2 81.57 34.78 13.08 34.25 35.45 36.00
- Co-rewarding-I (Ours) 67.6 83.01 32.22 13.50 32.38 35.56 35.53
- Co-rewarding-II (Ours) 66.2 80.89 33.28 14.40 32.88 36.94 37.59

### v3

@@ 4.2.1 Main Performance of Co-rewarding
4.2.2 Further Analysis
Co-rewarding alleviates collapse and provides stable self-supervised RL. We use 5,000 questions from the MATH test split as a validation set to monitor training process. Figure 4 shows that all three self-rewarding baselines collapse on both MATH and DAPO-14k. Co-rewarding-I remains stable on MATH but still collapses on DAPO-14k, suggesting that its stability depends on the property of training data. A plausible explanation is that the questions in MATH may provide favorable conditions for promoting diverse rephrasing variability, which is beneficial for the effectiveness of contrastive agreement in Co-rewarding-I. More discussions are provided in Appendix D.7. In contrast, Co-rewarding-II consistently maintains stability across datasets, as its design decouples supervision from the online policy and thus breaks the entanglement between supervision and the policy itself.
Each part contributes to Co-rewarding. Table 5 summarizes the ablations across two training sets. For Co-rewarding-I, we observe that it typically outperforms all three variants of Majority-Voting: models trained only on original questions, only on rephrased questions, or on their union. This indicates that the cross-supervision between original and rephrased questions plays a key role in mitigating training collapse, whereas simply adding more data does not resolve the inherent instability of single-view self-rewarding methods. Notably, training only on the original or rephrased data yields comparable results, reflecting that the quality of original and rephrased data is similar. For Co-rewarding-II, removing the EMA update of the reference teacher model causes clear degradation, highlighting the necessity of teacher updates for improving pseudo-label quality.
@@ 4.2.2 Further Analysis
Case Study of the model reasoning with different learning methods. Figure 7 provides a concrete example to illustrate the qualitative difference between self-rewarding baselines and our Co-rewarding. Majority-Voting exhibits reward hacking by boxing an incorrect answer “0” to pursue consensus, even though the reasoning steps are correct. Entropy produces repetitive outputs as its decoding probability distribution collapses onto a narrow set of tokens during entropy minimization. In contrast, Co-rewarding generates coherent step-by-step reasoning and correctly boxes the final answer, showing its capacity to overcome reward hacking and elicit genuine reasoning. Full results are provided in Appendix D.13 and additional case studies on code benchmark are discussed in Appendix D.14.
5 Related Work
Reinforcement learning with verifiable reward (RLVR) has recently become a mainstream post-training paradigm for eliciting strong reasoning abilities in LLMs (Guo et al., 2025), achieving remarkably encouraging success particularly on mathematical (Shao et al., 2024) and coding (Luo et al., 2025) tasks. However, RLVR fundamentally depends on high-quality and annotated GT labels to supervise reward signals, which remains a major bottleneck for scalability under the spirit of the scaling laws. To break this limitation, recent efforts have explored RL without external reward from multiple perspectives. For instance, methods such as TTRL (Zuo et al., 2025) and SRT (Shafayat et al., 2025) pursue self-consistency to generate pseudo labels for rewards, where agreement among multiple rollouts is treated as optimization objective. Additionally, another technical line such as EMPO (Zhang et al., 2025d), Intuitor (Zhao et al., 2025b) and RENT (Prabhudesai et al., 2025), enhances the LLM confidence by optimizing internal signals of reasoning, such as entropy minimization or self-certainty maximization. Different from these studies, Co-rewarding focuses on mitigating inherent training collapse in existing methods and enables stable self-supervised RL training. More detailed discussions of related work are in Appendix A.
@@ 6 Conclusion
In this work, we introduced Co-rewarding, a self-supervised RL framework that elicits the reasoning capability of LLMs through complementary supervision. Unlike prior self-rewarding methods that entangle rewards with single-view outputs and risk collapse, Co-rewarding establishes stability by decoupling the reward signal from the current online policy with the single-view output. Specifically, Co-rewarding-I leverages contrastive agreement across semantically analogous questions; Co-rewarding-II employs a dynamically updated teacher to provide insulated pseudo-labels; and Co-rewarding-III combines the data-side cross-supervision from Co-rewarding-I and the model-side teacher-based pseudo labels from Co-rewarding-II to further boost performance. Together, these designs construct cross-referable reward signals without explicit labels, aligning RL with invariances in reasoning rather than the mere correctness of isolated outputs. We hope this work will inspire further exploration into self-supervised RL for reasoning to advance the development.
Acknowledgments
ZZZ, JNZ, ZKZ, XL, XF and BH were supported by NSFC Major Research Plan No. 92570109.
Ethics Statement
This work complies with the Code of Ethics. It uses only publicly available datasets, involves no human or sensitive data, and raises no foreseeable risks related to privacy, security, or fairness issues. The research is conducted solely for scientific advancement, with no conflicts of interest.
Reproducibility Statement
We are committed to ensure the reproducibility of our proposed method. A detailed description of our approach is provided in the Co-rewarding Framework section, and the corresponding source code has been submitted in an anonymous repository at https://github.com/tmlr-group/Co-rewarding. Both backbone models and datasets used in our work are publicly available. Furthermore, all parameters, hyper-parameters, and procedural steps required to reproduce our results are thoroughly recorded in the Implementation Details. We believe that these components provide the community with details necessary to verify and reproduce our work.
@@ Appendix A Related Work
Large Language Model Reasoning. LLMs have shown impressive performance on vast tasks that require reasoning, including solving mathematical problems, writing code, and answering logical questions. One of the key techniques that has improved LLM reasoning is Chain-of-Thought (CoT) prompting (Wei et al., 2022; Zhou et al., 2024; 2025b). CoT encourages the model to generate intermediate reasoning steps before producing the final answer (Zhou et al., 2025c), which has been shown to enhance performance on tasks like arithmetic, commonsense reasoning, and symbolic reasoning. Subsequent work has extended CoT by integrating it with various strategies, including compositional generalization (Zhou et al., 2023a; Khot et al., 2023) and employing structural reasoning approaches (Yao et al., 2023a; Besta et al., 2024; Yang et al., 2024). In addition, CoT serves as a fundamental framework for techniques like fine-tuninig (Zelikman et al., 2022), argentic workflow (Yao et al., 2023b), and paving the way for inference-time scaling (Snell et al., 2024).
@@
RL for Large Language Models. Several RL algorithms have been developed primarily for alignment tasks (Wang et al., 2025b; Zhang et al., 2026; 2025a). Specifically, DPO (Rafailov et al., 2023), CPO (Xu et al., 2024), and their variants (Li et al., 2024b; Guo et al., 2024; Munos et al., 2024; Hong et al., 2024; Xie et al., 2024) rely on preference pairs labeled by annotators. In contrast, KTO (Ethayarajh et al., 2024) and BCO (Jung et al., 2024) require only a single binary label (like or dislike) for each output. Besides, PRM (Uesato et al., 2022; Lightman et al., 2024) and Step-KTO (Lin et al., 2025a) offer step-by-step guidance by incorporating feedback at each reasoning step rather than focusing solely on the final outputs. Recently, the follow-up work of GRPO improves the optimization objective, e.g., DAPO (Yu et al., 2025), Dr. GRPO (Liu et al., 2025a), REINFORCE++ (Hu, 2025), CPPO (Lin et al., 2025b), and GPG (Chu et al., 2025). Another line of research generalizes GRPO to broader applications such as multimodal reasoning (Zhou et al., 2025a; Huang et al., 2025; Chu et al., 2025; Liu et al., 2025b; Zhang et al., 2025b) and logical reasoning (Xie et al., 2025).
@@
RL without External Reward. RL methods have shown promising scaling capabilities to enhance the reasoning abilities of LLMs (Guo et al., 2025), yet they are often limited by the availability of training data for reward signals (Gao et al., 2023; Liu et al., 2023). Notably, Wang et al. (Wang et al., 2025a) demonstrate that RL can effectively bootstrap LLM reasoning with as little as a single training example, highlighting the potential to minimize or even eliminate reliance on external reward signals during training. Recent efforts leverage distinct strategies for reward assignment. For instance, SIRLC (Pang et al., 2024) and AZR (Zhao et al., 2025a) utilize an LLM-as-the-judge approach to assign rewards. In contrast, methods like SRT, TTRL, and their variants (Shafayat et al., 2025; Zuo et al., 2025; Fang et al., 2025; Zhang et al., 2025c) employ self-consistency (Wang et al., 2022) to generate pseudo-rewards, reducing dependence on external annotations. Meanwhile, INTUITOR, RLSC, and RENT (Zhao et al., 2025b; Li et al., 2025; Prabhudesai et al., 2025) harness the internal confidence scores of LLMs as intrinsic reward signals. Additionally, EMPO and its variants (Zhang et al., 2025d; Agarwal et al., 2025) promote reasoning by minimizing entropy of reasoning path, further diversifying the approaches to incentivize robust LLM performance (Han et al., 2025).
@@
Algorithm 3 Co-rewarding-III
1:Input: policy model $\pi_{\theta}$, learning rate $\eta$, original training dataset $\mathcal{D}$, rephrased training dataset $\mathcal{D^{\prime}}$, total iterations $K$.
2:Output: trained policy model $\pi_{\theta}$.
3:for iteration $k=1,\dots,K$ do
5:  for all $x\in\mathcal{B}$ and $x^{\prime}\in\mathcal{B}^{\prime}$ do
6:   Sample rollouts $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}^{(k)}(\cdot\mid x)$ and $\{y_{i}^{\prime}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}^{(k)}(\cdot\mid x^{\prime})$.
7:   Update the reference teacher by Eq. (11).
8:   Sample rollouts $\{\tilde{y}_{j}\}_{j=1}^{\tilde{G}}\sim\tilde{\pi}_{\mathrm{ref}}^{(k)}(\cdot\mid x)$ and $\{\tilde{y}_{j}^{\prime}\}_{j=1}^{\tilde{G}}\sim\tilde{\pi}_{\mathrm{ref}}^{(k)}(\cdot\mid x^{\prime})$.
9:   Obtain pseudo label from $\{\tilde{y}^{\prime}_{j}\}_{j=1}^{\tilde{G}}$ and $\{\tilde{y}_{j}\}_{j=1}^{\tilde{G}}$ by Eq. (13) and Eq. (14).
10:   Estimate the relative advantages by Eq. (13) and Eq. (14).
11:   Compute the objective by Eq. (3.2).
12:   Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{J}_{\mathrm{Co\text{-}rewarding\text{-}III}}^{(k)}(\theta)$.
13:  end for
14:end for
@@ Appendix B Pseudo Code of Co-rewarding
B.1 Formulation of Co-rewarding-III
The relative advantages $\mathcal{R}_{\theta}(\hat{A}^{(k)})$ and $\mathcal{R}_{\theta}(\hat{A}^{\prime(k)})$ are computed as:
$\hat{A}_{i}^{(k)}=\frac{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{\prime(k)}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})-\textrm{mean}(\{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{\prime(k)}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})\}_{i=1}^{G})}{\textrm{std}(\{r({\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{\prime(k)}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{i}})\}_{i=1}^{G})},\;{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}_{\textrm{v}}^{\prime(k)}}=\arg\max_{y*}\sum_{j=1}^{\tilde{G}}\mathbf{1}[\textrm{ans}({{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\tilde{y}^{\prime}_{j}}})=\textrm{ans}(y*)],$ (13)
$\hat{A}_{i}^{\prime(k)}=\frac{r({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{y}_{\textrm{v}}^{(k)}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y^{\prime}_{i}})-\textrm{mean}(\{r({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{y}_{\textrm{v}}^{(k)}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y^{\prime}_{i}})\}_{i=1}^{G})}{\textrm{std}(\{r({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{y}_{\textrm{v}}^{(k)}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}y^{\prime}_{i}})\}_{i=1}^{G})},\;{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{y}_{\textrm{v}}^{(k)}}=\arg\max_{y*}\sum_{j=1}^{\tilde{G}}\mathbf{1}[\textrm{ans}({{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{y}_{j}}})=\textrm{ans}(y*)],$ (14)
where the pseudo label $\tilde{y}_{\textrm{v}}^{\prime(k)}$ is the majority-vote pseudo label obtained from reference rollouts on the rephrased question, and $\tilde{y}_{\textrm{v}}^{(k)}$ is the corresponding pseudo label obtained from reference rollouts on original question. The reference model is slowly updated via EMA as in Eq. (11).
B.2 Pseudo Code
To intuitively present the pipeline of Co-rewarding, we summarize the pseudo codes of Co-rewarding-I, Co-rewarding-II and Co-rewarding-III in Algorithm 1, Algorithm 2 and Algorithm 3, respectively.
@@ 5. Output ONLY between the following markers, and strictly in this format (no extra explanation):
Settings Co-rewarding-I Co-rewarding-II Co-rewarding-III
Batch Size 128 128 128
Max Prompt Length 512 512 512
Max Response Length 3,072 3,072 3,072
Train Steps 170-220 (MATH), 300-330 (DAPO-14k), 100-130 (Open-RS)
Learning Rate 3e-6 3e-6 3e-6
# Policy Rollout $G$ 8 8 8
# Reference Rollout $\tilde{G} -$ 8 8
Clip Ratio 0.2 0.2 0.2
Warmup Style Cosine Cosine Cosine
Warmup Steps Ratio 0.1 0.1 0.1
KL Loss Coefficient 0.005 0.001 0.001
@@
Training Temperature 1.0 1.0 1.0
Evaluation Temperature 0.8 0.8 0.8
EMA $\alpha_{\textrm{start}}$ - 0.99 0.99
EMA $\alpha_{\textrm{end}}$ - 0.9999 0.9999
@@
Open-RS (Dang and Ngo, 2025) 7,000 Training Set
@@ C.3 More Evaluation Details
Methods MATH500 GSM8K AMC AIME24 LiveCode CRUX IFEval MMLU-Pro
@@
Before RL 53.6 19.48 10.69 0.52 9.95 18.50 29.83 32.50
- GT-Reward (Shao et al., 2024) 65.4 82.18 32.98 6.77 13.93 32.12 33.66 36.74
- Self-Certainty (Zhao et al., 2025b) 64.2 80.52 28.92 5.00 10.90 29.00 32.22 33.88
- Entropy (Prabhudesai et al., 2025) 63.2 80.44 29.67 5.94 9.05 29.00 32.94 35.35
- Majority-Voting (Shafayat et al., 2025) 64.6 82.41 33.13 5.10 14.03 36.38 35.19 35.50
- Co-rewarding-I (Ours) 65.4 84.53 30.57 5.31 16.40 36.88 33.86 36.38
- Co-rewarding-II (Ours) 65.2 81.72 32.38 4.47 22.25 40.25 32.74 30.79
@@
Before RL 69.4 24.71 15.81 2.81 3.79 26.38 38.19 44.76
- GT-Reward (Shao et al., 2024) 76.4 88.02 45.63 14.06 15.92 45.12 41.49 41.12
- Self-Certainty (Zhao et al., 2025b) 72.8 84.31 38.55 8.75 12.04 54.12 37.24 43.30
- Entropy (Prabhudesai et al., 2025) 72.2 81.43 39.61 10.73 16.49 51.88 40.33 42.79
- Majority-Voting (Shafayat et al., 2025) 74.4 84.53 40.96 11.04 15.45 51.00 38.60 43.35
- Co-rewarding-I (Ours) 74.6 89.61 41.27 10.73 15.73 55.58 42.86 40.51
- Co-rewarding-II (Ours) 73.6 89.31 42.77 11.98 8.25 47.50 41.82 37.45
@@
Before RL 57.0 19.56 8.43 1.15 4.45 7.50 33.65 33.00
- GT-Reward (Shao et al., 2024) 69.6 81.57 35.54 8.23 13.74 35.25 36.16 39.12
- Self-Certainty (Zhao et al., 2025b) 58.2 40.25 23.04 3.02 9.86 18.00 32.96 35.13
- Entropy (Prabhudesai et al., 2025) 63.6 71.79 31.63 6.88 13.74 31.37 35.37 36.67
- Majority-Voting (Shafayat et al., 2025) 65.2 81.57 34.78 7.50 13.08 34.25 35.45 36.00
- Co-rewarding-I (Ours) 67.6 83.01 32.22 8.65 13.50 32.38 35.56 35.53
- Co-rewarding-II (Ours) 66.2 80.89 33.28 7.50 14.40 32.88 36.94 37.59